OpenStudy (anonymous):

Solve the equation. five thriteenths t = –9

OpenStudy (anonymous):

(5/13)t=-9 t=-117/5

OpenStudy (anonymous):

\[(5/13)t =-9 ===> t=-9 *(13/5) = -117/5\]

OpenStudy (anonymous):

Thanks. :)